# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v2

@@ 2 Jack polynomial based hypergeometric function form
$\Big\langle(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}={s^{ql/2}\over l!B_{l}(q+1,\alpha)}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(q+l;(s)^{l})\Big|_{\alpha=1};$ (2.9)
@@
${1\over(2\pi)^{n}C_{\beta,n}}\prod_{1\leq j<k\leq n}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{\beta},\quad C_{\beta,n}={\Gamma(\beta n/2+1)\over\Gamma(\beta/2+1)^{n}};$ (2.10)
@@
$\Big\langle(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{{\rm CE}_{\beta,l}}={s^{ql/2}\over C_{\beta,l}B_{l}(q+1,2/\beta)}\,{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}(q+1+(l-1)\beta/2;(s)^{l}).$ (2.11)
For the $\beta$-generalisation of the eigenvalue PDF for complex Wishart matrices one has the functional form proportional to
$\prod_{l=1}^{N}x_{l}^{a}e^{-\beta x_{l}/2}\mathbbm{1}_{x_{l}>o}\prod_{1\leq j<k\leq N}|x_{k}-x_{j}|^{\beta};$ (2.12)
see e.g. [23, Eq. (3.16) with $a$ appropriately identified]. For this ensemble, let $E_{N,\beta}((0,s);a)$ denote the probability of there being no eigenvalues in $(0,s)$. Generalising the equality implied by (2.1) and (2.8), it was shown in [22] that for $a\in\mathbb{Z}_{\geq 0}$
$E_{\beta}^{\rm hard}((0,s);a):=\lim_{N\to\infty}E_{N,\beta}((0,s/(4N));a)=e^{-\beta s}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\beta/2)}(2a/\beta;(s/4)^{a}).$ (2.13)
@@ 3 Matrix differential equation characterisation
The significance of this is that for $\mathbf{x}=(t)^{n}$, since the 2022 work [28] we have available a matrix differential equation characterisation of ${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(a;b;\mathbf{x})$. This comes about in the setting of the integral representation [23, Cor. 2.1]
${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(2/\beta)}\Big((\beta/2)(n-1)+a+1;\beta(n-1)+a+b+2;(-s)^{n}\Big).\\ ={1\over S_{n}(a,b,\beta)}\int_{0}^{1}dx_{1}\cdots\int_{0}^{1}dx_{n}\,\prod_{l=1}^{n}x_{l}^{a}(1-x_{l})^{b}e^{-sx_{l}}\prod_{1\leq j<k\leq n}|x_{k}-x_{j}|^{\beta},$ (3.2)
@@
$H_{p}(x)={1\over C_{p}^{n}}\int_{0}^{1}dt_{1}\cdots\int_{0}^{1}dt_{n}\,\prod_{l=1}^{n}t_{l}^{a}(1-t_{l})^{b-1}e^{-x\sum_{l=1}^{n}t_{l}}\\ \times\prod_{1\leq j<k\leq n}|t_{k}-t_{j}|^{\beta}e_{p}(1-t_{1},\dots,1-t_{n}),$ (3.3)
for $p=0,1,\dots,n$. The factor $e_{p}(1-t_{1},\dots,1-t_{n})$ in the integrand denotes the $p$-th elementary symmetric polynomials in the variables $\{(1-t_{i})\}_{i=1}^{n}$, while $C_{p}^{n}$ is the binomial coefficient $n$ choose $p$. Setting $p=n$ gives the RHS of (3.2) and thus $_{1}F^{(2/\beta)}_{1}$, up to the Selberg integral normalisation, and setting $p=0$ gives the same with $b$ replaced by b1b-1.
Proposition 3.1.
[28, Cor. 2.5] The multidimesional integrals $\{H_{p}(x)\}_{p=0}^{n}$ satisfy the differential-difference system
$(n-p)xH_{p+1}(x)=\Big((n-p)x+B_{p}\Big)H_{p}(x)+x{d\over dx}H_{p}(x)-D_{p}H_{p-1}(x),\quad(p=0,\dots,n),$ (3.4)
@@
$B_{p}=p\Big(a+b+1+(\beta/2)(2n-p-1)\Big),\qquad D_{p}=p\Big((\beta/2)(n-p)+b\Big).$
@@
$\mathbf{X}$ $=-{\rm diag}\,[n,n-1,\dots,0]+{\rm diag}^{+}\,[n,n-1,\dots,1]$
$\mathbf{Y}$ $=-{\rm diag}\,[B_{0},B_{1},\dots,B_{n}]+{\rm diag}^{-}\,[D_{1},D_{2},\dots,D_{n}].$ (3.5)
Here $\rm diag$$^{+}$ ($\rm diag$$^{-}$) denotes the leading diagonal above (below) the diagonal itself. We have that the differential-difference equation (3.4) is equivalent to the matrix differential equation
$x{d\over dx}\mathbf{H}(x)=(x\mathbf{X}+\mathbf{Y})\mathbf{H}(x).$ (3.6)
@@
$\mathbf{H}(x)=\Big[\sum_{k=0}^{\infty}c_{p,k}x^{k}\Big]_{p=0}^{n},$
@@
$c_{0,0}=S_{n}(a,b-1,\beta),\quad c_{n,0}=S_{n}(a,b,\beta),\quad c_{n-p,0}=c_{n,0}\prod_{s=1}^{p}{B_{n+1-s}\over D_{n+1-s}},\>\>(p=1,\dots,n)$
@@
$(k{\mathbf{I}}_{n+1}-\mathbf{Y})\mathbf{c}_{k}=\mathbf{X}\mathbf{c}_{k-1},\>\>\mathbf{c}_{-1}:=\mathbf{0}\>\>(k=1,2,\dots),$ (3.7)
@@
Consider the differential-difference system [^1] Here and in the remainder of the paper the functions $\{I_{p}(x)\}$ are not to be confused with the I-Bessel function as appearing in (1.4) and (1.7)
$(n-p)xI_{p+1}(x)=B_{p}I_{p}(x)+x{d\over dx}I_{p}(x)-pI_{p-1}(x),\quad(p=0,\dots,n),$ (3.8)
@@
${\mathbf{W}}={\rm diag}^{+}\,[n,n-1,\dots,1],\quad{\mathbf{Z}}=-{\rm diag}\,[B_{0},B_{1},\dots,B_{n}]+{\rm diag}^{-}\,[1,2,\dots,n].$ (3.9)
@@
$x{d\over dx}\mathbf{I}(x)=(x\mathbf{W}+\mathbf{Z})\mathbf{I}(x).$ (3.10)
@@
$\mathbf{I}(x)=\Big[\sum_{k=0}^{\infty}d_{p,k}x^{k}\Big]_{p=0}^{n},$ (3.11)
@@
$I_{0}(x)$ $={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big(\beta(n-1)+a+b+1;(x)^{n}\Big),$ (3.12)
$I_{n}(x)$ $=d_{n,0}\,{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big(\beta(n-1)+a+b+2;(x)^{n}\Big),$ (3.13)
@@
$d_{p,0}={p!\over\prod_{s=1}^{p}B_{s}},\>\>(p=1,\dots,n).$ (3.14)
@@
$(k{\mathbf{I}}_{n+1}-\mathbf{Z})\mathbf{d}_{k}=\mathbf{W}\mathbf{d}_{k-1},\>\>(k=1,2,\dots).$ (3.15)
@@
$\mathbf{d}_{k}=(k{\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}\mathbf{W}\mathbf{d}_{k-1},\>\>(k=1,2,\dots)$ (3.17)
@@
1. Suppose we replace $q$ by q1q-1 in (3.16). Then setting $q=0$ we observe that the column vector $\delta_{p,l}]_{p=0}^{l}$ is an eigenvector of $\mathbf{Z}$} with $n=l$, allowing for the choice of initial vector corresponding to $d_{p,0}=\delta_{p,l}$ ($p=0,\dots,l$). Using this initial vector, an alternative formula to (3.18) is
$T_{l}(N)=(N!)^{2}(\mathbf{R}_{N,l}\,\mathbf{d}_{0})_{l+1},$ (3.19)
@@
2. For efficient computation using the matrix differential equation, it is advantageous to forgo (3.18) in favour of a recursive computation based on (3.15). See the second subsection of Appendix B for a complexity analysis when using this method to compute $\{T_{l}(N)\}$ in (1.3), which moreover is compared against the formalism used in [29, 12] for the same purpose.
@@ 4.1 The case of general parameters
Let $I_{n}(x)$ be specified via the differential-difference system (3.8). For some polynomials $\{q_{m}(x)\}_{m=0,1,\dots,n-1}$ we have
$\Big(x^{n}{d^{n+1}\over dx^{n+1}}+q_{n}(x)x^{n-1}{d^{n}\over dx^{n}}+q_{n-1}(x)x^{n-2}{d^{n-1}\over dx^{n-1}}+\cdots+q_{2}(x)x{d^{2}\over dx^{2}}+q_{1}(x){d\over dx}+q_{0}(x)\Big)I_{n}(x)=0.$ (4.1)
The degree in $x$ of the coefficient of each $d^{j}\over dx^{j}$} is [(n1+j)/2][(n-1+j)/2], thus restricting the degrees of the $q_{m}(x)$. Furthermore, when $\{B_{j}\}$ in (3.8) are all integers, the coefficients of these polynomials are also integers.
@@
Let [^2] This notation is not to be confused with $D_{p}$ in (3.4). $D_{x}:=x{d\over dx}$. From (3.8) in the case $p=n$ we read off that
@@
$n(n-1)(n-2)I_{n-3}(x)=\bigg((B_{n-2}+D_{x})\Big((B_{n-1}+D_{x})(B_{n}+D_{x})-nx\Big)-2x(n-1)(B_{n}+D_{x})\bigg)I_{n}(x).$ (4.4)
We see that the differential operators on the RHS of (4.2), (4.3), (4.4) have the structure
$x{d\over dx}+c_{0},\>\>x^{2}{d^{2}\over dx^{2}}+d_{0}x{d\over dx}+(c_{1}+c_{2}x),\>\>x^{3}{d^{3}\over dx^{3}}+e_{0}x^{2}{d^{2}\over dx^{2}}+(d_{1}+d_{2}x)x{d\over dx}+(c_{3}+c_{4}x),$ (4.5)
@@
${d\over dx}I_{0}(x)-nI_{1}(x)=0.$ (4.6)
From the established formulas for $I_{0}(x),I_{1}(x)$, this implies (4.1). That the coefficients of the polynomials $\{q_{j}(x)\}$ are integers for $\{B_{j}\}$ integers is immediate from the structure of the RHS of (4.2)–(4.4) as extends to $I_{0}(x),I_{1}(x)$. ∎
@@
In the case $\beta=2$, $n=l$ and $B_{p}=p(l-p)$, $I_{n}(x)$ is according to (2.9) and (3.13) (the latter modified in keeping with Remark 3.1.1) equal to the matrix average in (1.3) with $z=x^{1/2}$ and thus is a generating function for the sequence $\{T_{l}(N)\}_{N=1,2,\dots}$. Using different methods, the precise form of the corresponding differential operator in (4.1) has been computed for $l=2,3,\dots,7$ in [10]. From the substitution method given in the proof of Proposition 4.1 we can reclaim the listing of [10], and extend it to higher values of $l$ (see the work [45] for other applications in random matrix theory of this method in reducing first order linear matrix differential equations to higher order scalar differential equations). Here we make note of the case $l=8$.
@@
$q_{8}(x)=120,\>\>\>q_{7}(x)=12(483-10x),\>\>\>q_{6}(x)=4(36382-2541x),$
$q_{5}(x)=12(170417-27414x+364x^{2}),\>\>\>q_{4}(x)=24(214803+9830x),$
$q_{3}(x)=4(16531678-10189630x+1112127x^{2}-13120x^{3}),$
$q_{2}(x)=-8(-14762834+19053398x-4421047x^{2}+185712x^{3}),$
$q_{1}(x)=8(5883904-26926706x+13943655x^{2}-1483008x^{3}+18432x^{4}),$
$q_{0}(x)=256(-183872+380121x-98628x^{2}+4608x^{3}).$
In (1.3) let us replace $z$ by $x^{1/2}$ so that it reads
$\Big\langle e^{x^{1/2}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}x^{N}.$ (4.7)
As has just been illustrated in Proposition 4.2, we have from Proposition 4.1 that (4.7) satisfies a certain scalar linear differential equation of degree $l+1$. As remarked in this context in [10], it is well known and classical that one can translate the latter into a recurrence for $\{T_{l}(N)\}_{N=1,2,\dots}$ with coefficients that are polynomial in $N$. Previous literature contains the explicit form of the difference equations for small $l$; [12, footnote 22] gives references containing the explicit forms up to $l=5$. As an explicit example, from [52], for $l=4$ we have
$(N+4)(N+3)^{2}T_{4}(N)=(20N^{3}+62N^{2}+22N-24)T_{4}(N-1)-64N(N-1)^{2}T_{4}(N-2),$ (4.8)
valid for $N=1,2,\dots$. Iterating this, starting with $N=1,T_{4}(0)=1$ indeed generates the sequence (1.2). Below the proof of Proposition 4.4 note is made of the explicit forms for $l=2,3$ and 5.
@@
with each $P_{k}(N)$ a polynomial in $N$. However, already with $l=4$ we see from (4.8) that (4.9) is inaccurate, as (4.8) involves only $\{T_{4}(N),T_{4}(N-1),T_{2}(N-2)\}$ whereas (4.9) gives too a term involving $T_{2}(N-3)$. Making use of Proposition 4.1 we can (by inspection) give the corrected form of (4.9).
@@
Alternative to first determining the scalar differential equation and then converting it to a difference equation, it is possible to deduce the latter directly from the differential-difference system (3.8). [^3] We thank F. Bornemann for this observation.
@@
Substituting the power series for $I_{n-1}(x),I_{n}(x)$ in (4.2) and equating coefficients of $x^{N}$ gives
$d_{n-1,N}={1\over n}(B_{n}+N)d_{n,N}.$ (4.11)
Carrying out the analogous procedure in (3.8) with p=njp=n-j, $j=1,\dots,n-1$ gives
$d_{n-j-1,N}={1\over n-j}\Big((B_{n-j}+N)d_{n-j,N}-jd_{n-j+1,N-1}\Big).$ (4.12)
After successive substitutions, at the completion of step $j$, one has an expression for $d_{n-j-1,N}$ in terms of $\{d_{n,N^{\prime}}\}_{N^{\prime}=N-1-[j/2],\dots,N}$. In relation to $d_{0,N}$ and $d_{1,N}$ there is the additional equation
$(N+1)d_{0,N+1}-nd_{1,N}=0$ (4.13)
as follows from (4.6). Substituting the two previously obtained expressions for $d_{0,N},d_{1,N}$ in terms of $\{d_{n,N^{\prime}}\}$ we deduce (4.10). ∎
As noted in the first line of this subsection, in the present context, $n=l,B_{p}=p(l-p)$. Further, with $\{T_{l}(N)\}$ as in (4.10) we have $d_{l,N}=T_{l}(N)/(N!)^{2}$. In terms of $\{T_{l}(N)\}$, in low order cases the procedure of the proof of Proposition 4.4 gives
$(N+2)T_{2}(N+1)=(4N+2)T_{2}(N),\quad(N+3)^{2}T_{3}(N+1)=(9+22N+10N^{2})T_{3}(N)-9N^{2}T_{3}(N-1).$ (4.14)

### v3

@@ 2 Jack polynomial based hypergeometric function form
$\big\langle(\det U)^{q}{\rm e}^{s^{1/2}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}={s^{ql/2}\over l!B_{l}(q+1,\alpha)}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}\bigl(q+l;(s)^{l}\bigr)\big|_{\alpha=1};$ (2.5)
@@
${1\over(2\pi)^{n}C_{\beta,n}}\prod_{1\leq j<k\leq n}\big|{\rm e}^{{\rm i}\theta_{k}}-{\rm e}^{{\rm i}\theta_{j}}\big|^{\beta},\qquad C_{\beta,n}={\Gamma(\beta n/2+1)\over\Gamma(\beta/2+1)^{n}};$ (2.6)
@@
$\big\langle(\det U)^{q}{\rm e}^{s^{1/2}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{{\rm CE}_{\beta,l}}={s^{ql/2}\over C_{\beta,l}B_{l}(q+1,2/\beta)}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\bigl(q+1+(l-1)\beta/2;(s)^{l}\bigr).$ (2.7)
For the $\beta$-generalisation of the eigenvalue PDF for complex Wishart matrices, one has the functional form proportional to
$\prod_{l=1}^{N}x_{l}^{a}{\rm e}^{-\beta x_{l}/2}\mathbbm{1}_{x_{l}>o}\prod_{1\leq j<k\leq N}|x_{k}-x_{j}|^{\beta};$
see, for example, [24, equation (3.16) with $a$ appropriately identified]. For this ensemble, let $E_{N,\beta}((0,s);a)$ denote the probability of there being no eigenvalues in $(0,s)$. Generalising the equality implied by (2.1) and (2.4), it was shown in [23] that for $a\in\mathbb{Z}_{\geq 0}$
$E_{\beta}^{\rm hard}((0,s);a):=\lim_{N\to\infty}E_{N,\beta}((0,s/(4N));a)={\rm e}^{-\beta s}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\beta/2)}(2a/\beta;(s/4)^{a}).$
@@ 3 Matrix differential equation characterisation
The significance of this is that for $\mathbf{x}=(t)^{n}$, since the 2022 work [28] we have available a matrix differential equation characterisation of ${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(a;b;\mathbf{x})$. This comes about in the setting of the integral representation [24, Corollary 2.1]
${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(2/\beta)}((\beta/2)(n-1)+a+1;\beta(n-1)+a+b+2;(-s)^{n}).$
$\qquad={1\over S_{n}(a,b,\beta)}\int_{0}^{1}{\rm d}x_{1}\cdots\int_{0}^{1}{\rm d}x_{n}\prod_{l=1}^{n}x_{l}^{a}(1-x_{l})^{b}{\rm e}^{-sx_{l}}\prod_{1\leq j<k\leq n}|x_{k}-x_{j}|^{\beta},$ (3.2)
@@
$H_{p}(x)={1\over C_{p}^{n}}\int_{0}^{1}{\rm d}t_{1}\cdots\int_{0}^{1}{\rm d}t_{n}\prod_{l=1}^{n}t_{l}^{a}(1-t_{l})^{b-1}{\rm e}^{-x\sum_{l=1}^{n}t_{l}}$
[…]
for $p=0,1,\dots,n$. The factor $e_{p}(1-t_{1},\dots,1-t_{n})$ in the integrand denotes the $p$-th elementary symmetric polynomials in the variables $\{(1-t_{i})\}_{i=1}^{n}$, while $C_{p}^{n}$ is the binomial coefficient $n$ choose $p$. Setting $p=n$ gives the right-hand side of (3.2) and thus $_{1}F^{(2/\beta)}_{1}$, up to the Selberg integral normalisation, and setting $p=0$ gives the same with $b$ replaced by b1b-1.
Proposition 3.1 ([28, Corollary 2.5]).
The multi-dimensional integrals $\{H_{p}(x)\}_{p=0}^{n}$ satisfy the differential-difference system
$(n-p)xH_{p+1}(x)=((n-p)x+B_{p})H_{p}(x)+x{{\rm d}\over{\rm d}x}H_{p}(x)-D_{p}H_{p-1}(x),$
$p=0,\dots,n,$ (3.3)
@@
$B_{p}=p(a+b+1+(\beta/2)(2n-p-1)),\qquad D_{p}=p((\beta/2)(n-p)+b).$
@@
$\mathbf{X}$ $=-\operatorname{diag}[n,n-1,\dots,0]+\operatorname{diag}^{+}[n,n-1,\dots,1],$
$\mathbf{Y}$ $=-\operatorname{diag}[B_{0},B_{1},\dots,B_{n}]+\operatorname{diag}^{-}[D_{1},D_{2},\dots,D_{n}].$
Here $\operatorname{diag}^{+}$ ($\operatorname{diag}^{-}$) denotes the leading diagonal above ((below)) the diagonal itself. We have that the differential-difference equation (3.3) is equivalent to the matrix differential equation
$x{{\rm d}\over{\rm d}x}\mathbf{H}(x)=(x\mathbf{X}+\mathbf{Y})\mathbf{H}(x).$ (3.4)
@@
$\mathbf{H}(x)=\left[\sum_{k=0}^{\infty}c_{p,k}x^{k}\right]_{p=0}^{n},$
@@
$c_{0,0}=S_{n}(a,b-1,\beta),\qquad c_{n,0}=S_{n}(a,b,\beta),\qquad c_{n-p,0}=c_{n,0}\prod_{s=1}^{p}{B_{n+1-s}\over D_{n+1-s}},$
$p=1,\dots,n$
@@
$(k{\mathbf{I}}_{n+1}-\mathbf{Y})\mathbf{c}_{k}=\mathbf{X}\mathbf{c}_{k-1},\qquad\mathbf{c}_{-1}:=\mathbf{0},\qquad k=1,2,\dots,$ (3.5)
@@
Consider the differential-difference system [^1] Here and in the remainder of the paper the functions $\{I_{p}(x)\}$ are not to be confused with the I-Bessel function as appearing in (1.4) and (1.7).
$(n-p)xI_{p+1}(x)=B_{p}I_{p}(x)+x{{\rm d}\over{\rm d}x}I_{p}(x)-pI_{p-1}(x),\qquad p=0,\dots,n,$ (3.6)
@@
${\mathbf{W}}=\operatorname{diag}^{+}[n,n-1,\dots,1],\qquad{\mathbf{Z}}=-\operatorname{diag}[B_{0},B_{1},\dots,B_{n}]+\operatorname{diag}^{-}[1,2,\dots,n].$
@@
$x{{\rm d}\over{\rm d}x}\mathbf{I}(x)=(x\mathbf{W}+\mathbf{Z})\mathbf{I}(x).$ (3.7)
@@
$\mathbf{I}(x)=\left[\sum_{k=0}^{\infty}d_{p,k}x^{k}\right]_{p=0}^{n},$ (3.8)
@@
$I_{0}(x)$ $={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}(\beta(n-1)+a+b+1;(x)^{n}),$ (3.9)
$I_{n}(x)$ $=d_{n,0}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}(\beta(n-1)+a+b+2;(x)^{n}),$ (3.10)
@@
$d_{p,0}={p!\over\prod_{s=1}^{p}B_{s}},\qquad p=1,\dots,n.$ (3.11)
@@
$(k{\mathbf{I}}_{n+1}-\mathbf{Z})\mathbf{d}_{k}=\mathbf{W}\mathbf{d}_{k-1},\qquad k=1,2,\dots.$ (3.12)
@@
$\mathbf{d}_{k}=(k{\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}\mathbf{W}\mathbf{d}_{k-1},\qquad k=1,2,\dots$ (3.14)
@@
  • (1)
    Suppose we replace $q$ by q1q-1 in (3.13). Then setting $q=0$ we observe that the column vector $\delta_{p,l}]_{p=0}^{l}$ is an eigenvector of $\mathbf{Z}$} with $n=l$, allowing for the choice of initial vector corresponding to $d_{p,0}=\delta_{p,l}$ ($p=0,\dots,l$). Using this initial vector, an alternative formula to (3.15) is
    $T_{l}(N)=(N!)^{2}(\mathbf{R}_{N,l}\mathbf{d}_{0})_{l+1},$
@@
  • (2)
    For efficient computation using the matrix differential equation, it is advantageous to forgo (3.15) in favour of a recursive computation based on (3.12). See Appendix B.2 for a complexity analysis when using this method to compute $\{T_{l}(N)\}$ in (1.3), which moreover is compared against the formalism used in [13, 29] for the same purpose.
@@ 4.1 The case of general parameters
Let $I_{n}(x)$ be specified via the differential-difference system (3.6). For some polynomials $\{q_{m}(x)\}_{m=0,1,\dots,n-1}$, we have
$\left(x^{n}{{\rm d}^{n+1}\over{\rm d}x^{n+1}}+q_{n}(x)x^{n-1}{{\rm d}^{n}\over{\rm d}x^{n}}+q_{n-1}(x)x^{n-2}{{\rm d}^{n-1}\over{\rm d}x^{n-1}}+\cdots+q_{2}(x)x{{\rm d}^{2}\over{\rm d}x^{2}}+q_{1}(x){{\rm d}\over{\rm d}x}+q_{0}(x)\right)$
$\qquad\times I_{n}(x)=0.$ (4.1)
The degree in $x$ of the coefficient of each ${\rm d}^{j}\over{\rm d}x^{j}$} is [(n1+j)/2][(n-1+j)/2], thus restricting the degrees of the $q_{m}(x)$. Furthermore, when $\{B_{j}\}$ in (3.6) are all integers, the coefficients of these polynomials are also integers.
@@
Let [^2] This notation is not to be confused with $D_{p}$ in (3.3). $D_{x}:=x{{\rm d}\over{\rm d}x}$. From (3.6) in the case $p=n$, we read off that
@@
$n(n-1)(n-2)I_{n-3}(x)$
$\qquad=((B_{n-2}+D_{x})((B_{n-1}+D_{x})(B_{n}+D_{x})-nx)-2x(n-1)(B_{n}+D_{x}))I_{n}(x).$ (4.4)
We see that the differential operators on the right-hand sides of (4.2), (4.3), (4.4) have the structure
$x{{\rm d}\over{\rm d}x}+c_{0},\qquad x^{2}{{\rm d}^{2}\over{\rm d}x^{2}}+d_{0}x{{\rm d}\over{\rm d}x}+(c_{1}+c_{2}x),$
$x^{3}{{\rm d}^{3}\over{\rm d}x^{3}}+e_{0}x^{2}{{\rm d}^{2}\over{\rm d}x^{2}}+(d_{1}+d_{2}x)x{{\rm d}\over{\rm d}x}+(c_{3}+c_{4}x),$
@@
${{\rm d}\over{\rm d}x}I_{0}(x)-nI_{1}(x)=0.$ (4.5)
From the established formulas for $I_{0}(x)$, $I_{1}(x)$, this implies (4.1). That the coefficients of the polynomials $\{q_{j}(x)\}$ are integers for $\{B_{j}\}$ integers is immediate from the structure of the right-hand sides of (4.2)–(4.4) as extends to $I_{0}(x)$, $I_{1}(x)$. ∎
@@
In the case $\beta=2$, $n=l$ and $B_{p}=p(l-p)$, $I_{n}(x)$ is according to (2.5) and (3.10) (the latter modified in keeping with Remark 3.4.1) equal to the matrix average in (1.3) with $z=x^{1/2}$ and thus is a generating function for the sequence $\{T_{l}(N)\}_{N=1,2,\dots}$. Using different methods, the precise form of the corresponding differential operator in (4.1) has been computed for $l=2,3,\dots,7$ in [11]. From the substitution method given in the proof of Proposition 4.1, we can reclaim the listing of [11], and extend it to higher values of $l$ (see the work [45] for other applications in random matrix theory of this method in reducing first-order linear matrix differential equations to higher-order scalar differential equations). Here we make note of the case $l=8$.
@@
$q_{8}(x)=120,\qquad q_{7}(x)=12(483-10x),\qquad q_{6}(x)=4(36382-2541x),$
$q_{5}(x)=12\bigl(170417-27414x+364x^{2}\bigr),\qquad q_{4}(x)=24(214803+9830x),$
$q_{3}(x)=4\bigl(16531678-10189630x+1112127x^{2}-13120x^{3}\bigr),$
$q_{2}(x)=-8\bigl(-14762834+19053398x-4421047x^{2}+185712x^{3}\bigr),$
$q_{1}(x)=8\bigl(5883904-26926706x+13943655x^{2}-1483008x^{3}+18432x^{4}\bigr),$
$q_{0}(x)=256\bigl(-183872+380121x-98628x^{2}+4608x^{3}\bigr).$
In (1.3), let us replace $z$ by $x^{1/2}$ so that it reads
$\big\langle{\rm e}^{x^{1/2}\operatorname{Tr}(U+U^{\dagger})}\big\rangle_{U(l)}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}x^{N}.$ (4.6)
As has just been illustrated in Proposition 4.2, we have from Proposition 4.1 that (4.6) satisfies a certain scalar linear differential equation of degree $l+1$. As remarked in this context in [11], it is well known and classical that one can translate the latter into a recurrence for $\{T_{l}(N)\}_{N=1,2,\dots}$ with coefficients that are polynomial in $N$. Previous literature contains the explicit form of the difference equations for small $l$; [13, footnote 22] gives references containing the explicit forms up to $l=5$. As an explicit example, from [52], for $l=4$ we have
$(N+4)(N+3)^{2}T_{4}(N)={}$ $\bigl(20N^{3}+62N^{2}+22N-24\bigr)T_{4}(N-1)$
$-64N(N-1)^{2}T_{4}(N-2),$ (4.7)
valid for $N=1,2,\dots$. Iterating this, starting with $N=1,T_{4}(0)=1$ indeed generates the sequence (1.2). Following the proof of Proposition 4.4 below, note is made of the explicit forms for $l=2,3$ and 5.
@@
with each $P_{k}(N)$ a polynomial in $N$. However, already with $l=4$ we see from (4.7) that (4.8) is inaccurate, as (4.7) involves only $\{T_{4}(N),T_{4}(N-1),T_{2}(N-2)\}$ whereas (4.8) gives too a term involving $T_{2}(N-3)$. Making use of Proposition 4.1, we can (by inspection) give the corrected form of (4.8).
@@
Alternative to first determining the scalar differential equation and then converting it to a difference equation, it is possible to deduce the latter directly from the differential-difference system (3.6). [^3] We thank Folkmar Bornemann for this observation.
@@
Substituting the power series for $I_{n-1}(x),I_{n}(x)$ in (4.2) and equating coefficients of $x^{N}$ gives $d_{n-1,N}={1\over n}(B_{n}+N)d_{n,N}$. Carrying out the analogous procedure in (3.6) with p=njp=n-j, $j=1,\dots,n-1$ gives
$d_{n-j-1,N}={1\over n-j}((B_{n-j}+N)d_{n-j,N}-jd_{n-j+1,N-1}).$
After successive substitutions, at the completion of step $j$, one has an expression for $d_{n-j-1,N}$ in terms of $\{d_{n,N^{\prime}}\}_{N^{\prime}=N-1-[j/2],\dots,N}$. In relation to $d_{0,N}$ and $d_{1,N}$ there is the additional equation $(N+1)d_{0,N+1}-nd_{1,N}=0$ as follows from (4.5). Substituting the two previously obtained expressions for $d_{0,N}$, $d_{1,N}$ in terms of $\{d_{n,N^{\prime}}\}$, we deduce (4.9). ∎
As noted in the first line of this subsection, in the present context, $n=l,B_{p}=p(l-p)$. Further, with $\{T_{l}(N)\}$ as in (4.9) we have $d_{l,N}=T_{l}(N)/(N!)^{2}$. In terms of $\{T_{l}(N)\}$, in low-order cases the procedure of the proof of Proposition 4.4 gives
$(N+2)T_{2}(N+1)=(4N+2)T_{2}(N),$
$(N+3)^{2}T_{3}(N+1)=\bigl(9+22N+10N^{2}\bigr)T_{3}(N)-9N^{2}T_{3}(N-1).$ (4.10)